Project

General

Profile

Actions

2017-07-20-LH2

6th July 2017; 11:00 BST

KL, CW, MT, AK, JB

Phone details

Notes

Actions

  • Action AK follow through de-bugging of FC datalogger. - Trigger signal not in archiver - signal triggers at 3.3V, was at 8V, now reduced a system is off - requires wire trace - Action on CMcW.
  • Action CMcW to install NI on MICEnet machine. Complete
  • Action MT/JB to agree and implement testing of quench valves. IN pogress
  • Action PW to modify purge sequence limits in agreement with JB. In Progress
  • Action AK/PW to determine required connections HMI to PLC. Complete - requires PW testing of permissions.. Action PW/AK.
    • Action CMcW/JB to search archiver for FC radiation shield temp & helium pressure excursions. Ongoing
  • Action PF/AK to propose IOC test program for next week. New expert panel has been tested - system details from PLC are not reported(can be removed) but information retrieved is correct. some additions have been agreed - implement tomorrow.
  • Action AK/JB to arrange further meeting on alarm handler development. Fri 21st July
  • Action JB/CW/MT to review SH report. Complete - single 100mm line is sufficient
  • Action CW to retrieve RP notes on testing on H2 panel. - checked notes were with Kieran. MT now repeating all tests to allow completing required test documentation.
  • Action CW update project plan. complete, attached
  • Action CW to contact AN to review wiki and paperwork. Superceded JB/PF to agree wiki page.

Preliminary H2 tour

Actions recorded by AN. Post today.
nitrogen purge of vent line, reviewers noted this should be purged with nitrogen to prevent hydrogen venting into a line filled with air. Check of full P&ID shown this is in place.
atex rating of 'eveything' in H2 shed.
vacuum pump motors, heaters - certificates? Action JB.

external components - feedthroughs piercing wall. Seal with foam.

note final review of paperwork is required - requested MICE/ISIS LH2 meeting next week to generate complete list of required paperwork. Acton KL to contact JT

Monitoring and archival of data - PF/AK.

discuss possible additinal archived variables beyond those for monitoring

Current technical

  • FC pressure rating.
  • Refurbishment of pipework from turret to wall.
  • Refurbishment of pipework from wall to roof and vent.

Neon Run

purge and cool Thursday/Friday.

Shifters

  • Atex training Additonal staff for Atex on 20th 21st Sept.
    JB to discuss with DH H2 experts without Atex.

Hall access

Timetable

  • Next meeting - Thurs 20th July
  • AoB

H2 rota based on 'stable system' requiring only on-call care. On duty period may extend over 24hrs. Action JB
Action on CW to contact HR re 'on-call' times.

SW - Kl contacted GR - requires reply limited request on SW time.

O Kirichek magnet meeting pre-meeting early August.

Back to LH2:

Back

Files:

Updated by Whyte, Colin almost 6 years ago · 2 revisions